OpenStudy (anonymous):

Find the number of real number solution of 4x^2+16x+15=0?

OpenStudy (jackellyn):

Factor 4x^2+10x+6x+15 = 0 2x(2x+5) + 3(2x+5) = 0 (2x+5)(2x+3) = 0 x = -5/2 or x = -3/2

OpenStudy (jiteshmeghwal9):

\[4x^2+10x+6x+15=0\]\[=>2x(2x+5)+3(2x+5)\]\[(2x+3)(2x+5)\]

OpenStudy (jiteshmeghwal9):

\[x=\frac{-5}{2} or \frac{-3}{2}\]

OpenStudy (jiteshmeghwal9):

Is that clear now???